# Moonshine: Speech Recognition for Live Transcription and Voice Commands

### v1

@@ 1 Introduction
Our first task was to better quantify the bottleneck imposed by Whisper’s fixed-length encoder. In an empirical test, we compared the GFLOPS (i.e., billion floating-point operations) needed for processing 30 seconds of zero-padded audio to the GFLOPS that would be required to instead process only a <30absent30<30< 30 second subset of the same audio—i.e., we measured the potential for speed-ups if Whisper instead had a variable-length encoder. Our results (Figure 1) showed potential for a 35x speed-up in the best case, and nearly a 5x speed-up overall. Clearly, improvements to Whisper’s architecture could yield lightweight, low-latency speech-to-text models that are better-suited to resource-constrained applications.
@@ 3.2 Training data collection & preprocessing
Preprocessing noisily-labeled speech. Many speech sources available on the web have subtitles or captions available, which can serve as labels. However, captions tend be noisy—they may be manually-generated and thus contain text that is orthogonal to the audio content, or they may contain the names of speakers or verbal descriptions of non-speech content. In cases where a manually-generated but possibly-unreliable caption is available, we use a heuristic process to filter out low-quality instances. First, we lowercase and normalize the caption text, removing or replacing, e.g., ambiguous unicode characters, emoji, and punctuation. We then use Whisper large v3 to generate a pseudo-label of the audio content, applying the same text normalization to this pseudo-label as we do the caption. Finally, we compute a normalized Levenshtein distance (between [0.0,1.0]0.01.0[0.0,1.0][ 0.0 , 1.0 ], where $0.00.0$ is identical and $1.01.0$ is orthogonal) between the normalized caption and the pseudo-label, filtering out labels with a distance above a threshold. This allows us to treat the human-generated labels in captions as ground truth without introducing excessive noise. After filtering out noisy labels, we prepare the remaining text by applying standardized punctuation and capitalization.
@@
[…]

### v2

@@ 1 Introduction
Our first task was to better quantify the bottleneck imposed by Whisper’s fixed-length encoder. In an empirical test, we compared the GFLOPS (i.e., billion floating-point operations) needed for processing 30 seconds of zero-padded audio to the GFLOPS that would be required to instead process only a <30<30 second subset of the same audio—i.e., we measured the potential for speed-ups if Whisper instead had a variable-length encoder. Our results (Figure 1) showed potential for a 35x speed-up in the best case, and nearly a 5x speed-up overall. Clearly, improvements to Whisper’s architecture could yield lightweight, low-latency speech-to-text models that are better-suited to resource-constrained applications.
@@ 3.2 Training data collection & preprocessing
Preprocessing noisily-labeled speech. Many speech sources available on the web have subtitles or captions available, which can serve as labels. However, captions tend be noisy—they may be manually-generated and thus contain text that is orthogonal to the audio content, or they may contain the names of speakers or verbal descriptions of non-speech content. In cases where a manually-generated but possibly-unreliable caption is available, we use a heuristic process to filter out low-quality instances. First, we lowercase and normalize the caption text, removing or replacing, e.g., ambiguous unicode characters, emoji, and punctuation. We then use Whisper large v3 to generate a pseudo-label of the audio content, applying the same text normalization to this pseudo-label as we do the caption. Finally, we compute a normalized Levenshtein distance (between $[0.0,1.0]$, where $0.0$ is identical and $1.0$ is orthogonal) between the normalized caption and the pseudo-label, filtering out labels with a distance above a threshold. This allows us to treat the human-generated labels in captions as ground truth without introducing excessive noise. After filtering out noisy labels, we prepare the remaining text by applying standardized punctuation and capitalization.
@@
Controlling instance duration. We assemble successive speech segments into longer training instances, such that the resulting instance duration is $\in[4,30]$ seconds with no more than $2$ seconds between successive segments. For our manually-captioned data sources, we use timestamped segments provided by a subtitle file (e.g., an .srt file); for our pseudo-labeled audio, we use the timestamps output by Whisper. Upon combining this data with open datasets, the duration of instances in our aggregate training set obeys a slightly bimodal distribution as depicted in Figure 4.